OpenStudy (anonymous):

Find the exact value of cos165 degrees

OpenStudy (anonymous):

hi what are you choices

OpenStudy (anonymous):

I don't have any lol

OpenStudy (anonymous):

ok

OpenStudy (sbuck98):

Here use this cosine calcualtor http://rapidtables.com/calc/math/Cos_Calculator.htm

OpenStudy (anonymous):

do you mean the cosine of 165 degrees?

OpenStudy (anonymous):

Then -(√ (3) + 1)√ (2) / 4

OpenStudy (anonymous):

ohhhh. okay thank you. and I cant use a calc but thanks

OpenStudy (jdoe0001):

got Unit circle?

OpenStudy (mathmale):

Find the exact value of cos165 degrees Have you considered using the addition formula for the cosine function? Note that cos 75 = cos (45 + 30) = cos 45 cos 30 - sin 45 sin 30. Think about this. How could you use this juicy tidbit to solve the problem at hand?

OpenStudy (jdoe0001):

\(\bf 165=120+45\qquad thus\qquad cos(165^o)\iff cos(120^o+45^o) \\ \quad \\ recall \implies cos({\color{brown}{ \alpha}} + {\color{blue}{ \beta}})= cos({\color{brown}{ \alpha}})cos({\color{blue}{ \beta}})- sin({\color{brown}{ \alpha}})sin({\color{blue}{ \beta}}) \\ \quad \\ cos({\color{brown}{ 120^o}})cos({\color{blue}{ 45^o}})-sin({\color{brown}{ 120^o}})sin({\color{blue}{ 45^o}})\)

OpenStudy (anonymous):

yeah thanks guys I understand. the book gives me a formula but I'm dumb and didnt check. Thanks though

OpenStudy (mathmale):

This is an excellent approach. Note that both 120 and 45 degrees are "special angles." What's special about them?

OpenStudy (anonymous):

They have values from the unit circle... i guess